Little Dome C - Beyond EPICA Drilling Site (75.29917 degS, 122.44516 degE)

Situation Report #10, 11. December 2021

Personnel @LDC: Saverio Panichi (ENEA, Camp Leader), Carlo Barbante (ISP-CNR, PI), Olivier Alemany (IGE), Olivier Delanoe (IPEV), Romain Duphil (IPEV), Matthias Hüther (AWI), Calogero Monaco (ENEA), Anthony Pauty (IPEV), Philippe Possenti (IPEV), Michele Scalet (ENEA), Jakob Schwander (unibe), Thomas Stocker (unibe), Gregory Teste (IGE), Remo Walther (unibe)

Personnel @DC: Giacomo Bonanno (ENEA)

Weather at LDC 9 pm: sunny

Meteo at DC 9 pm: T = −29°C, Wind = 3 knt, Wind Chill T = −35°C

Reaming continued in the BEOIC drilling tent. The second step started today according to the schedule. The two Vissmann cabins were connected to electricity. Also the AWI workshop received proper power. Michele covered the core buffer containers partially by moving many cubic meters of snow with the PB 300. This will guarantee stable cold temperatures for the BEOIC ice cores.

BEOIC drilling depth at the end of the day: 128.0 m

Calogero and Greg took a ride to Dome Concordia transporting back recycling material. Also, Greg exchanged with Giacomo Bunoanno, the satellite communication specialist. Calogero and Giacomo returned in time for the Saturday evening dinner, including the famous LDC Apero.

Giacomo is eagerly awaited to put the new broadband system into operation. Too many Whatsapp messages and emails are in the pipeline by now. We apologize to our family and friends that this type of communication has been scarce until now.

RADIX firn drilling progressed for a few meters but encountered some problems due to the stiff nylon tube which exerted too much friction along the borehole. The 120 meter tube was heated and stretched so that it became straight. We leave the tube in the stretched position overnight and continue with the firn drilling tomorrow.

Radix drilling depth: 22 m

Thomas Stocker and Carlo Barbante